OpenStudy (anonymous):

How do you factor the following polymonial completely; 245v^2-180?

OpenStudy (uri):

245v^2-180 5(49v^2-34)

OpenStudy (goformit100):

for factoring these type of question, first try to solve it using 'Theory of equations.

OpenStudy (agent0smith):

First factor out the GCF, then see if it's a difference of two squares, and continue factoring if it is.

OpenStudy (agent0smith):

@uri, 5*34 is not equal to 180 :P

OpenStudy (agent0smith):

Factor out the GCF of 5, \[\Large 245v^2-180 = 5(49v^2 - 36)\] Notice what's inside the brackets is a difference of two squares and can be written like so \[\Large 5(49v^2 - 36) = 5 \left[ (7v)^2 - (6)^2 \right]\] Diff of two squares is: \[\Large a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

Thx, agent0smith. I appreciate ur help with this...